OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5

OpenStudy (tommo_lcfc):

The best thing to do is to use substitution to answer this question. You can re-arrange the second equation to get \[x = -5 - 2y = -(5+2y)\] Then substitute this into the first equation to get: \[2(-5-2y) + 3y = 4\] This can be simplified to \[-10 - 7y = 4\] so 7y = 14 and y = 2. Substitute this into the second equation (this is easiest!) to get x - 4 = -5, so x = -1 Summarising, x = -1 and y = 2. Hope this helps!

OpenStudy (anonymous):

WOW! Thanks so much!